OpenStudy (anonymous):

Can anyone show the work for using the difference quitent for x^3

OpenStudy (anonymous):

\[f(x)=x^3, \frac{f(a+h)-f(a)}{h}\]

OpenStudy (anonymous):

(f(a+h)-f(a))/h =((a+h)^3-a^3)/h =(h^3+3h^2a+3ha^2+a^3-a^3)/h =(h^3+3h^2a+3ha^2)/h =h(h^2+3ha+3a^2)/h =h^2+3ha+3a^2 its usually used when h->0 so =3a^2

OpenStudy (anonymous):

*lim h->0

OpenStudy (anonymous):

I came up with 2a^2+2ah+h^2 So now i'm attempting to figure where i went wrong, thanks for the help

OpenStudy (anonymous):

np